OpenStudy (anonymous):

The effective rate of a $25,000 non-interest-bearing simple discount 10%, 90-day note is?

OpenStudy (anonymous):

help

OpenStudy (tkhunny):

Payback: 25000.00 10% Discount: 25000.00*0.10*(90/365) = 616.44 <== You may need to use 360. You didn't say. Pocket Money: 25000.00 - 616.44 = 24383.56 Effective Rate: 616.44 = 24383.56*r*(90/365) -- Solve for r Do it again using 360 in place of 365. See how different the answer is.